Jo is X years old. Jo's mother , Anne, is four times as old as Jo. In four years time , anne will be three times as old as Jo. Form an equation in X and solve it to find jo's present age
Answer by ankor@dixie-net.com(22740)

Jo is X years old. Jo's mother,
x = Joe's age
:
Anne, is four times as old as Jo.
4x = Anne's age
:
In four years time, Anne will be three times as old as Jo.
4x + 4 = 3(x + 4)
:
Form an equation in X and solve it to find jo's present age
4x + 4 = 3(x+4)
4x + 4 = 3x + 12
4x - 3x = 12 - 4
x = 8 yrs is Joe's age